Density is a measure of the mass per unit of volume for a material. Mathematically, this can be expressed in units of kg/m3. Given that the equation for the volume of sphere is based on radius, the density of Earth is ___x103 kg/m3



1
Expert's answer
2021-11-21T17:28:00-0500

Let's first find the volume of Earth:


"V=\dfrac{4}{3}\pi r^3=\dfrac{4}{3}\pi\times(6.37\times10^6\ m)^3=1.08\times10^{21}\ m^3."

Finally, we can find the density of Earth:


"\rho=\dfrac{m}{V}=\dfrac{5.98\times10^{24}\ kg}{1.08\times10^{21}\ m^3}=5.537\times10^3\ \dfrac{kg}{m^3}."